OpenStudy (australia10):

how do i visualise 0.03 / 0.1 3 hundredeths divided by 1 tenth I understand that It is just just 3/100 x 10 as you get the reciprocal of 1/10 but I struggle to understand how 0.1 divides in .03 visually.

OpenStudy (australia10):

.1 doesn't go into 0.03 unless you get the recirical of 1/10 which is 10 and times it bhy 0.03.

ganeshie8 (ganeshie8):

Nice question. May be lets look at an example problem : Imagine you have a `bag full of chocolates` and you want to distribute them to `students in a classroom`.

OpenStudy (australia10):

If it was 0.3/0.1 it would be easy enough but if its 0.03/0.1 i get confused.

ganeshie8 (ganeshie8):

How is it easy if it were 0.3/0.1 ?

OpenStudy (australia10):

.1 goes into .3 = 3 times

ganeshie8 (ganeshie8):

Do you mean 0.1 x 3 = 0.3 ?

OpenStudy (australia10):

I don't think so. .3/.1=3 as .3 x 10=3

OpenStudy (australia10):

but 0.1 doesnt go into 0.03

OpenStudy (australia10):

Not sure if I ake any sense sorry.

ganeshie8 (ganeshie8):

You're absolutely correct. 0.1 doesn't go into 0.03 directly. For one reason, 0.1 is greater than 0.03. How can you divide something so small (0.03) by something so big (0.1) ? Doesn't look plausible. But we don't stop here. There will be situations where we want to perform this kind of division anyways. Let's look at the example problem I gave you earlier..

ganeshie8 (ganeshie8):

Also assume that your bag has thousands of of chocolates, but you want to give away only 3/100 th of them.

ganeshie8 (ganeshie8):

Does the below phrase make sense ? "3/100 th of the chocolates in your bag"

OpenStudy (australia10):

but 0.03/.1 = .3 so arent we still dividing it somehow?

OpenStudy (australia10):

That makes sense.

ganeshie8 (ganeshie8):

Yes, we are able to divide; its just that the result 0.3 is not a good looking integer

ganeshie8 (ganeshie8):

Good. Let me ask a quick question. If you have 1000 chocolates in your bag, how many of them you're going to give away ?

OpenStudy (australia10):

30/1000 = 3/100

OpenStudy (australia10):

? i think

ganeshie8 (ganeshie8):

Nope. Please read my question again. I think if the question makes sense, you will be able to answer it correctly :)

ganeshie8 (ganeshie8):

What is 3/100th of 1000 ?

ganeshie8 (ganeshie8):

If you have 100 chocolates, you will be giving away 3 chocolates. If you have 1000 chocolates, how many will you be giving away ?

OpenStudy (australia10):

wouldnt the equivalent in thousandths be 30 chocolates? because you x10 to both numerator and demoninator?

ganeshie8 (ganeshie8):

Yes!

ganeshie8 (ganeshie8):

Here is the story so far : You have a lot of chocolates in you bag and you want to give away 3/100 of them to the students.

OpenStudy (australia10):

would that make the equation 0.0030 / 0.1 ?

ganeshie8 (ganeshie8):

But wait, you don't really want to give the chocolates to all the students. You just want to give the chocolates to only 1/10 th of the students.

ganeshie8 (ganeshie8):

Again, a quick question : If there are 50 students in the class, how many of them will be getting your chocolates ?

OpenStudy (australia10):

1/10 of them?

ganeshie8 (ganeshie8):

Yes, what's 1/10 th of 50 ?

OpenStudy (australia10):

5

ganeshie8 (ganeshie8):

Excellent! Thank you for surviving so far and responding brilliantly :) Let's keep going..

ganeshie8 (ganeshie8):

Here is the story so far : You have a lot of chocolates in you bag and you want to give away 3/100 of them to the students. Also you want to give the chocolates to only 1/10th of the students.

OpenStudy (australia10):

Yep that makes sense.

OpenStudy (australia10):

so far

ganeshie8 (ganeshie8):

Since that situation makes sense, below expression also should make sense. \[\dfrac{3/100\times \color{blue}{1000}}{1/10\times \color{blue}{50}}\]

ganeshie8 (ganeshie8):

\(\color{blue}{1000}\) chocolates are there in your bag. \(\color{blue}{50}\) students are there in the class.

ganeshie8 (ganeshie8):

The expression in the top : \(3/100\times \color{blue}{1000}\) represents below : 3/100th of the chocolates in your bag.

ganeshie8 (ganeshie8):

similarly the expression in the bottom : \(1/10\times \color{blue}{50}\) represents below : 1/10th of the students in the class.

ganeshie8 (ganeshie8):

still with me ?

OpenStudy (australia10):

im not sure how you get the 50 students

OpenStudy (australia10):

I have to go for today. But I appreciate the help.

OpenStudy (australia10):

I've been thinking could I say it's 3 of the chocolates devided by 0.1 30 but 30 is x100 overvalued so its 30/100 which simplifies to 0.1

OpenStudy (australia10):

i mean 0.3

OpenStudy (australia10):

?

OpenStudy (mathmale):

Starting with 0.03 / 0.1 with the goal of making the arithmetic easier: eliminate the decimal fractions. You could accomplish this by multiplying both numerator and denom. by 100. @australia10, would you please do this and share your result here.

OpenStudy (phi):

how do i visualise 0.03 / 0.1 the same way you would "visualise" 3/10 If you can make sense of 3/10 , then use that same idea to understand 0.03/.1 (it's the same fraction, but scaled down by 100)